I got this only math homework and I really need help solving it. This is what it says: Bryan bought a new battery-gasoline hybrid car. On a trip the car ran exclusively on its battery for the first 40 miles, and then ran exclusively on gasoline, using gasoline at a rate of .02 gallons per mile. On the whole trip he averaged 60 miles per gallon. How long was the trip in miles?
Answer by Edwin McCravy(20077)

I got this only math homework and I really need help solving it. This is what it says: Bryan bought a new battery-gasoline hybrid car. On a trip the car ran exclusively on its battery for the first 40 miles, and then ran exclusively on gasoline, using gasoline at a rate of .02 gallons per mile. On the whole trip he averaged 60 miles per gallon. How long was the trip in miles?

Let x = the answer, the number of miles long the trip was.
Then x-40 = the number of miles the car went on gasoline.
Then .02(x-40) = the number of gallons used on the whole trip

     x miles            60 miles
-----------------  =  ------------ 
.02(x-40) gallons       1 gallon

Cross multiply

      60(.02)(x-40) = x
      
          1.2(x-40) = x

           1.2x-48 = x

               .2x = 48

                 x = 240

240 miles.

Edwin